OpenStudy (anonymous):

2√(50)/5√(2) show work please.

OpenStudy (shamil98):

\[\large (12*7^{1/2})^2 \] \[\large 12^2 * 7^{1/2 *2} = 12^2 * 7^1\] Can you simplify it from here?

OpenStudy (shamil98):

You can also use this rule.. \[\large \sqrt{x}^2 = \sqrt{x} * \sqrt{x} = x\]

OpenStudy (shamil98):

\[\large \sqrt{x} = x^{1/2}\] if you're wondering where the 1/2 came from.

OpenStudy (anonymous):

when I simplify it, I got the answer?

OpenStudy (shamil98):

Yes.

OpenStudy (anonymous):

@shamil98 2√(50)/5√(2) ?

OpenStudy (shamil98):

No.. What is (12)^2 * 7 ?

OpenStudy (shamil98):

Unless you are asking a different questions

OpenStudy (anonymous):

the answer to my first question is 1008. my second question is 2√(50)/5√(2)

OpenStudy (anonymous):

@shamil98 the answer with my first question is 1008, correct? my second question is 2√(50)/5√(2)

OpenStudy (shamil98):

You're correct on the first one.

OpenStudy (shamil98):

\[\large \frac{ 2 \sqrt{50} }{ 5\sqrt{2} }\] \[\huge \frac{ 2 \sqrt{25*2} }{ 2\sqrt{5}} = \frac{ \cancel 2*5 \sqrt{2} }{ \cancel 2\sqrt{5} }\]

OpenStudy (shamil98):

\[\huge \frac{ 5\sqrt{2} }{ \sqrt{5} } \times \frac{ \sqrt{5} }{ \sqrt{5} } = \frac{ \cancel 5\sqrt{10} }{ \cancel 5 } = \sqrt{10}\]

OpenStudy (anonymous):

@shamil98 are you sure the answer isn't 2?

OpenStudy (shamil98):

yes..

OpenStudy (anonymous):

okay my third question is (11√(5)(7√(10)

OpenStudy (anonymous):

@shamil98 my third question is (11√(5)(7√(10)

OpenStudy (anonymous):

@shamil98 √(1200)/√(588) ?